# RIGNO: A Graph-based framework for robust and accurate operator learning for PDEs on arbitrary domains

### v1

@@ Acknowledgements
The contribution of S. Mishra to this work was supported in part by the DOE SEA-CROGS project (DE-SC-0023191).
Supplementary Material for RIGNO: A Graph-based Framework
for robust and accurate operator learning for PDEs on arbitrary domains.
@@ A.4 Fixed and Variable physical meshes
(a) Support sub-regions of regional nodes.
@@
Appendix B Boundary Conditions in RIGNO
B.1 Periodic boundary conditions
@@
In order to motivate the above changes, let us consider two nodes, one close to the left and one close to the right boundary. With (normalized) raw coordinates, the initial features of these two nodes will be very different as one will take a value close to 00 and the other one takes a value close to $1$. With periodic boundary conditions, however, these two points will be close to each other if the frame of the reference domain is slightly shifted. The invariance to a shift in the frame of reference means that these two nodes should be connected in the graphs. In simpler words, we assume during graph construction that the both ends of a domain with periodic boundary conditions are connected. In one-dimensional settings, this set-up corresponds to considering a circular domain, see Figure B.1 for an illustration. For a square two-dimensional domain, connecting the two ends in both direction corresponds to a torus.
@@
Inspired by the ghost node technique in classical numerical methods, we extend the domain in all directions by constructing a 3$\times$3 tiled structure, where each tile contains the repeated coordinates of the given point cloud. All the Delaunay triangulations in the graph construction phase are built based on the extended domain. Figure A.1 shows the resulting support sub-regions, as well as the extra edges that are introduced because of the domain extensions for a minimal set of regional nodes. We refer to these extra edges as cross-boundary edges. Notice how all the white areas in Figure 1(a) are covered by at least on circle on the other side of the domain. In order to avoid introducing extra costs to the architecture, the ghost nodes are discarded after graph construction and only the cross-boundary edges are kept. In order to keep the extra edges, we replace each ghost node with its corresponding node in the original point cloud. The result is illustrated in Figure 1(d).
@@
Figure B.1: Graph structure for periodic boundary conditions in one-dimensional settings. The green dots represent physical nodes, the purple circles represent regional nodes, and the solid lines represent edge connections. Note that only a few physical nodes are visualized. The bottom right part of the figure partly shows the receptive fields of two regional nodes in the encoder. Here, the red lines show the edges that are randomly masked, and hence will not be used in that particular message passing block.
B.2 General boundary conditions
In problems with general (non-periodic) boundary conditions, we rely on the capacity of the model to learn the boundary conditions implicitly from data. Although this seems to be sufficient for the problems considered in this work, we would like to highlight that more sophisticated mechanisms are needed for non-homogeneous boundary conditions or homogeneous boundary conditions that change in each sample. One possible approach is to consider a separate type of edge or node for encoding data on the boundaries, similar to the type extension method that is proposed in [32], to differentiate between mesh edges and proximity-based edges for Lagrangian systems. In this work, however, we only consider one node type and one edge type, since the considered solution operators are not dominated by boundary conditions.
@@ C.2 Temporal Inference
In all the time-dependent datasets considered in this work, for simplicity of exposition, we have considered $N_{t}=20$ or $N_{t}=21$ time snapshots $\{t_{n}\}_{n=1}^{N_{t}}$. In all trainings, only snapshots with even indices up to $t_{14}$ are used. Snapshots with $t>t_{14}$ are reserved for time extrapolation tests, and snapshots with $t<t_{14}$ with even indices are reserved for time interpolation tests. Error at snapshot $t_{14}$ is used as a reference for the benchmarks. Since the trained operator is continuous in time and can be inferred with different lead times, there are infinitely many ways to reach the solution at $t_{14}$ via autoregressive inference of the trained operator. The optimal autoregressive path can be different for different operators. Figure C.2 illustrates the training and test time snapshots as well as different autoregressive paths for reaching $t_{14}$. In the presented benchmarks in Table 1, we calculate the error with 3 different autoregressive paths: i) $\{\tau_{i}=2\Delta t\}_{i=1}^{7}$ which is denoted by AR-2; ii) $\{\tau_{1}=4\Delta t,\tau_{2}=4\Delta t,\tau_{3}=4\Delta t,\tau_{4}=2\Delta t\}$ which is denoted by AR-4; and iii) $\{\tau_{1}=14\Delta t\}$, denoted by DR. We always report the lowest error among these three rollout paths.
@@ C.3 Fractional Pairing
We use the fractional pairing strategy to achieve a truly time-continuous operator that can be inferred with lead times smaller than the smallest time difference in the training dataset. In fractional pairing, we approximate the solution at an intermediary time by inferring a model that has been trained on the original time resolution. This approximation is then given as input to the same model. By choosing a suitable intermediary time, we can have an input-output pair that is not present in the training dataset and have a smaller lead time than the time resolution of the training dataset. This idea is illustrated in Figure C.3. With this, we can either calculate the loss gradients for both of the inferences, or we can cut the gradient at the intermediary time and treat the intermediary solution as a normal model input, which is referred to as GradCut (see bottom left of Figure C.3). In the experiments of this work, we always use fractional pairing with GradCut.
If the approximation at the intermediary time is not accurate enough, fractional pairing strategy can lead to training instabilities. Therefore, it is recommended to use this technique only as a fine-tuning stage. Furthermore, since the approximations are made with fractional lead times, i.e., lead times that are not present in the training dataset, it is also essential to employ other mechanisms that help generalize to fractional lead times with regular training strategies, such as consistent time stepping strategies and time-conditioned normalization layers. A further consideration that helps the stability of this strategy is avoiding smaller-than-time-resolution lead times in the first forward pass. Such small lead times can be seen as an extrapolation in the seen distribution of lead times during training. It has been observed that even without fractional pairing, the models generalize better to interpolated lead times than they do to extrapolated lead times. Therefore, we can approximate the intermediary solution with an interpolated lead time, and use a smaller lead time only in the second forward pass. In Figure C.3, this corresponds to only using the two right-most additional pairs in the green box. Note that none of the additional pairs of the orange box will be viable in this setting.
Figure C.3: Schematic of the fractional pairing strategy. The additional pairs obtained by fractional pairing correspond to smaller lead times than the smallest time difference in the original trajectory. On the bottom left of the figure, the forward passes (solid black arrows) and the backward gradient flows (dashed arrows) are illustrated for an existing pair $(t_{m},t_{n})$. With the GradCut setting, the gradient flow is stopped at $t_{\ast}$.
@@ Appendix D Datasets
RIGNO is tested on a suite of 12 datasets [31] which covers operators arising from compressible Euler (CE) equations of gas dynamics, incompressible Navier-Stokes (NS) equations of fluid dynamics, wave equation (WE), heat equation (HE) of thermal dynamics, Allen-Cahn equation (ACE), and force balance equation in solid mechanics (FB). Table D.1 summarizes the main characteristics of all the datasets. A random sample of each dataset is visualized in Figures D.2-D.13. This collection of datasets covers diverse multi-scale, nonlinear, complex phenomena in systems arising from PDEs such as shock waves, shear layers, mixing, and turbulence. This makes the effort to learn the underlying solution operator from data worthwhile for these datasets; since running simulations of these systems with any classical numerical method would require extensive computational resources. Moreover, by considering these datasets, we cover a diverse set of domains and meshes from Cartesian uniform grids on a unit square to structured and unstructured meshes on fixed and varying irregular geometries.
The problems in the 8 datasets from [12] are all solved in the unit spatio-temporal domain $\Omega_{t}\times\Omega_{x}=[0,T]\times[0,1]^{2}$. The final time is $T=1$ except for the ACE dataset where $T=0.0002$. For each dataset, a parameterized distribution of initial conditions is considered, and initial conditions are randomly drawn by sampling those distributions. The distribution of the initial conditions characterizes the dataset, the level of complexity of the dynamics captured in it, the phenomenological features exhibited in the solution function, and the corresponding solution operator. In all these datasets, $N_{t}=21$ snapshots of each sample are collected at times $\Omega_{t}^{\Delta}=\{t_{n}=n\Delta t\}_{n=0}^{N_{t}-1}$ with uniform spacing $\Delta t$; except for the ACE and the Wave-Layer datasets where $N_{t}=20$. Each snapshot originally has a 128$\times$128 resolution, with uniform discretization in both spatial directions. In order to test the baselines on a random point cloud with these datasets, we randomly shuffle the coordinates and only keep the first 9216 coordinates in our experiments. Figure D.1 illustrates a few random point clouds obtained from this procedure.
Figure D.1: Random point clouds (9216 points) obtained from a 128$\times$128 uniform grid.
@@
Abbreviation PDE BC Domain Mesh Source Visualization
Heat-L-Sines HE Dirichlet L-shape unstructured generated Figure D.2
Wave-C-Sines WE Dirichlet circle unstructured generated Figure D.3
NS-Gauss INS periodic square uniform grid [12] Figure D.4
NS-PwC INS periodic square uniform grid [12] Figure D.5
NS-SL INS periodic square uniform grid [12] Figure D.6
NS-SVS INS periodic square uniform grid [12] Figure D.7
CE-Gauss CE periodic square uniform grid [12] Figure D.8
CE-RP CE periodic square uniform grid [12] Figure D.9
ACE ACE periodic square uniform grid [12] Figure D.10
Wave-Layer WE absorbing square uniform grid [12] Figure D.11
AF CE various airfoil structured [21] Figure D.12
Elasticity FB various holed square unstructured [21] Figure D.13
@@ D.1.1 Incompressible Navier-Stokes equations (INS)
[…]
@@ D.1.5 Heat equation (HE)
D.1.6 Solid body force balance equation (FB)
@@ D.2.1 Heat-L-Sines
This dataset considers time-dependent solutions of the heat equation over an L-shaped domain which is formed by removing the rectangular portion [1, 1.5]×[0.5, 1]11.50.51[1,\,1.5$\times$[0.5,\,1][ 1 , 1.5 ] × [ 0.5 , 1 ] from the larger rectangle [0.5, 1.5]×[0, 1]0.51.501[0.5,\,1.5$\times$[0,\,1][ 0.5 , 1.5 ] × [ 0 , 1 ]. We consider constant diffusivity $a=1.0$ and zero homogeneous Dirichlet boundary conditions. We prescribe a superposition of sinusoidal modes as the initial condition:
@@
where $\boldsymbol{\mu}\in[-1,1]^{d}$, $M=10$, and each $\mu_{m}$ is sampled independently and uniformly from $[-1,1]$. A Delaunay-based triangular mesh with characteristic length $0.008$ is used, yielding about $14{,}047$ nodes and $27{,}590$ elements. We simulate the heat equation for $100$ time steps with a time step size of $\Delta t=5\times 10^{-5}$ and keep $21$ (including the initial condition) equally spaced snapshots from each trajectory.

### v2

@@ Acknowledgements
The contribution of Siddhartha Mishra to this work was supported in part by the DOE SEA-CROGS project (DE-SC-0023191).
Supplementary Material
RIGNO: A Graph-based Framework for Robust and Accurate
Operator Learning for PDEs on Arbitrary Domains
@@ A.4 Fixed and variable physical meshes
(a) Aggregation area of regional nodes.
@@
A.5 Periodic boundary conditions
@@
In order to motivate the above changes, let us consider two nodes, one close to the left and one close to the right boundary. With (normalized) raw coordinates, the initial features of these two nodes will be very different as one will take a value close to 0 and the other one takes a value close to $1$. With periodic boundary conditions, however, these two points will be close to each other if the frame of the reference domain is slightly shifted. The invariance to a shift in the frame of reference means that these two nodes should be connected in the graphs. In simpler words, we assume during graph construction that the both ends of a domain with periodic boundary conditions are connected. In one-dimensional settings, this set-up corresponds to considering a circular domain, see Figure A.2 for an illustration. For a square two-dimensional domain, connecting the two ends in both direction corresponds to a torus.
@@
Inspired by the ghost node technique in classical numerical methods, we extend the domain in all directions by constructing a $\times3$ tiled structure, where each tile contains the repeated coordinates of the given point cloud. All the Delaunay triangulations in the graph construction phase are built based on the extended domain. Figure A.1 shows the resulting support sub-regions, as well as the extra edges that are introduced because of the domain extensions for a minimal set of regional nodes. We refer to these extra edges as cross-boundary edges. Notice how all the white areas in Figure 1(a) are covered by at least on circle on the other side of the domain. In order to avoid introducing extra costs to the architecture, the ghost nodes are discarded after graph construction and only the cross-boundary edges are kept. In order to keep the extra edges, we replace each ghost node with its corresponding node in the original point cloud. The result is illustrated in Figure 1(d).
@@
Figure A.2: Graph structure for periodic boundary conditions in one-dimensional settings. The green dots represent physical nodes, the purple circles represent regional nodes, and the solid lines represent edge connections. Note that only a few physical nodes are visualized. The bottom right part of the figure partly shows the receptive fields of two regional nodes in the encoder. Here, the red lines show the edges that are randomly masked, and hence will not be used in that particular message passing block.
A.6 Structural features
The corresponding coordinates of the nodes are pre-processed to provide the initial structural features of the graph entities. In order to impose translation equivariance to the model, we consider linear normalized coordinates $\boldsymbol{\zeta}_{i}\in[-1,+1]$ and angular normalized coordinates $\boldsymbol{\alpha}_{i}\in[0,2\pi]$ defined as $\boldsymbol{\zeta}_{i}:=2{(\boldsymbol{\rm x}_{i}-\boldsymbol{\rm x}_{\rm c})}/{\boldsymbol{L}}-1$ and $\boldsymbol{\alpha}_{i}:=\pi(\boldsymbol{\zeta}_{i}+1)$. The structural features of an edge $\boldsymbol{e}_{ij}$ connecting node $i$ to node $j$ is $\boldsymbol{\varphi}_{ij}:=[\tilde{\boldsymbol{\zeta}}_{ij}\diamond\|\tilde{\boldsymbol{\zeta}}_{ij}\|_{2}]$; with $\tilde{\boldsymbol{\zeta}}_{ij}:=(\boldsymbol{\zeta}_{j}-\boldsymbol{\zeta}_{i})/(2\sqrt{d})$ being the relative coordinates of the connected nodes normalized by the diameter of the domain, $2\sqrt{d}$. We denote the vertical concatenation of column vectors by $\diamond$. The structural features of a node $\boldsymbol{v}_{i}$ with general boundary conditions are simply its normalized coordinates $\boldsymbol{\varphi}_{i}:=\boldsymbol{\zeta}_{i}$. With periodic boundary conditions (see Section A.5), we concatenate sines and cosines of angular normalized coordinates with $K_{\rm freq}$ frequencies: $\boldsymbol{\varphi}_{i}=\left[\sin(k\boldsymbol{\alpha}_{i})\diamond\cos(k\boldsymbol{\alpha}_{i})\right]$ for $1\leq k\leq K_{\rm freq}$.
Appendix B Details of the GNN Components
We consider two underlying sets of nodes (see Figure 1). $\boldsymbol{V}_{\mathfrak{P}}=\{\boldsymbol{v}_{i}^{\mathfrak{P}}\}_{i}$ denotes the physical nodes, which corresponds to the unstructured point cloud on which the input (and output) data is defined. The regional nodes are denoted by $\boldsymbol{V}_{\mathfrak{R}}=\{\boldsymbol{v}_{i}^{\mathfrak{R}}\}_{i}$ and the bidirectional edges that connect them by $\boldsymbol{E}_{\mathfrak{R}\leftrightarrow\mathfrak{R}}=\{\boldsymbol{e}_{ij}^{\mathfrak{R}\leftrightarrow\mathfrak{R}}\}_{ij}$. $\boldsymbol{E}_{\mathfrak{P}\to\mathfrak{R}}=\{\boldsymbol{e}_{ij}^{\mathfrak{P}\to\mathfrak{R}}\}_{ij}$ and $\boldsymbol{E}_{\mathfrak{R}\to\mathfrak{P}}=\{\boldsymbol{e}_{ij}^{\mathfrak{R}\to\mathfrak{P}}\}_{ij}$ denote the unidirectional edges in the encoder and the deocder, respectively.
The first step is to embed the initial features of all nodes and all edges (see Section A.6) to suitable high-dimensional features. The other model inputs (initial condition, spatial parameters, and time) are only embedded in the features of the physical nodes. For regional nodes, we also embed the radius of their support sub-region $R_{i}$. This is done by passing all the initial features through the embedder feed-forward blocks (FF, see Figure 1),
$\boldsymbol{v}^{\mathfrak{P},(0)}_{i}$ $=\textsl{FF}_{\text{embedder}}^{\mathfrak{P}}(\boldsymbol{u}^{t}_{i},\boldsymbol{\varphi}_{i},\boldsymbol{c}^{t}_{i},t,\tau)$ $\emph{\small(Physical nodes)},$ (5)
$\boldsymbol{v}^{\mathfrak{R},(0)}_{i}$ $=\textsl{FF}_{\text{embedder}}^{\mathfrak{R}}(\boldsymbol{\varphi}_{i},R_{i})$ $\emph{\small(Regional nodes)},$
$\boldsymbol{e}^{\mathfrak{P}\to\mathfrak{R},(0)}_{ij}$ $=\textsl{FF}_{\text{embedder}}^{\mathfrak{P}\to\mathfrak{R}}(\boldsymbol{\varphi}_{ij})$ $\emph{\small(Physical to regional edges)},$
$\boldsymbol{e}^{\mathfrak{R}\leftrightarrow\mathfrak{R},(0)}_{ij}$ $=\textsl{FF}_{\text{embedder}}^{\mathfrak{R}\leftrightarrow\mathfrak{R}}(\boldsymbol{\varphi}_{ij})$ $\emph{\small(Regional mesh edges)},$
$\boldsymbol{e}^{\mathfrak{R}\to\mathfrak{P},(0)}_{ij}$ $=\textsl{FF}_{\text{embedder}}^{\mathfrak{R}\to\mathfrak{P}}(\boldsymbol{\varphi}_{ij})$ $\emph{\small(Regional to physical edges)}.$
Next, the encoder applies a single message passing step to the embedded features of the physical nodes and transmits the information to the regional mesh,
The processor passes this latent information on the regional mesh through $P$ sequential message-passing blocks with residual updates and skip connections. The $p$-th message-passing step in the processor updates the regional nodes and the regional mesh edges as,
The decoder then transfers processed information from the regional mesh back to the physical mesh by performing a single message passing step,
Finally, in the second step of the decoder, the feature vectors of the physical node are independently scaled down to the desired dimension to form the output
@@ C.2 Temporal inference
In all the time-dependent datasets considered in this work, for simplicity of exposition, we have considered $N_{t}=20$ or $N_{t}=21$ time snapshots $\{t_{n}\}_{n=1}^{N_{t}}$. Unless otherwise stated, only snapshots with even indices up to $t_{14}$ are used for training. Snapshots with $t>t_{14}$ are reserved for time extrapolation tests, and snapshots with $t<t_{14}$ with even indices are reserved for time interpolation tests. Error at snapshot $t_{14}$ is used as a reference for the benchmarks. Since the trained operator is continuous in time and can be inferred with different lead times, there are infinitely many ways to reach the solution at $t_{14}$ via autoregressive inference of the trained operator. The optimal autoregressive path can be different for different operators. Figure C.2 illustrates the training and test time snapshots as well as different autoregressive paths for reaching $t_{14}$. In the presented benchmarks in Tables 1 and 2, we calculate the error with 3 different autoregressive paths: i) $\{\tau_{i}=2\Delta t\}_{i=1}^{7}$ which is denoted by AR-2; ii) $\{\tau_{1}=4\Delta t,\tau_{2}=4\Delta t,\tau_{3}=4\Delta t,\tau_{4}=2\Delta t\}$ which is denoted by AR-4; and iii) $\{\tau_{1}=14\Delta t\}$, denoted by DR. We always report the lowest error among these three rollout paths.
@@ C.3 Fractional pairing
We use the fractional pairing strategy to achieve a truly time-continuous operator that can be inferred with lead times smaller than the smallest time difference present in the training dataset. In fractional pairing, we approximate the solution at an intermediary time by inferring a model that has been trained on the original time resolution. This approximation is then given as input to the same model. By choosing a suitable intermediary time, we can have an input-output pair that is not present in the training dataset and have a smaller lead time than the time resolution of the training dataset. This idea is illustrated in Figure C.3. With this, we can either calculate the loss gradients for both of the inferences, or we can cut the gradient at the intermediary time and treat the intermediary solution as a normal model input, which is referred to as GradCut (see bottom left of Figure C.3). In the experiments of this work, we always use fractional pairing with GradCut.
If the approximation at the intermediary time is not accurate enough, fractional pairing strategy can lead to training instabilities. Therefore, it is recommended to use this technique only as a fine-tuning stage. Furthermore, since the approximations are made with fractional lead times, i.e., lead times that are not present in the training dataset, it is also essential to employ other mechanisms that help generalize to fractional lead times with regular training strategies, such as consistent time marching strategies and time-conditioned normalization layers. A further consideration that helps the stability of this strategy is avoiding out-of-distribution lead times in the first forward pass, i.e., avoiding lead times that are smaller than the time resolution of the training dataset. Such small lead times can be seen as an extrapolation in the seen distribution of lead times during training. It has been observed that even without fractional pairing, the models generalize better to interpolated lead times than they do to extrapolated ones. Therefore, we can approximate the intermediary solution with an interpolated lead time, and use a smaller lead time only in the second forward pass. In Figure C.3, this corresponds to only using the two right-most additional pairs in the green box. Note that none of the additional pairs of the orange box will be viable in this setting.
Figure C.3: Schematic of the fractional pairing strategy. The additional pairs obtained by fractional pairing correspond to smaller lead times than the smallest time difference in the original trajectory. On the bottom left of the figure, the forward passes (solid black arrows) and the backward gradient flows (dashed arrows) are illustrated for an existing pair $(t_{m},t_{n})$. With the GradCut setting, the gradient flow is stopped at $t_{\star}$.
@@ Appendix D Datasets
RIGNO is tested on a suite of 13 datasets [32] which covers operators arising from compressible Euler (CE) equations of gas dynamics, incompressible Navier-Stokes (NS) equations of fluid dynamics, wave equation (WE), heat equation (HE) of thermal dynamics, Allen-Cahn equation (ACE), Poisson’s equation, and force balance equation in solid mechanics (FB). Table D.1 summarizes the main characteristics of all the datasets. A random sample of each dataset is visualized in Figures D.2-D.14. This collection of datasets covers diverse multi-scale, nonlinear, complex phenomena in systems arising from PDEs such as shock waves, shear layers, mixing, and turbulence. This makes the effort to learn the underlying solution operator from data worthwhile for these datasets; since running simulations of these systems with any classical numerical method would require extensive computational resources. Moreover, by considering these datasets, we cover a diverse set of domains and meshes from Cartesian uniform grids on a unit square to structured and unstructured meshes on fixed and varying irregular geometries.
The problems in the 8 time-dependent datasets from [13] are all solved in the unit spatio-temporal domain $\Omega_{t}\times\Omega_{x}=[0,T]\times[0,1]^{2}$. The final time is $T=1$ except for the ACE dataset where $T=0.0002$. For each dataset, a parameterized distribution of initial conditions is considered, and initial conditions are randomly drawn by sampling those distributions. The distribution of the initial conditions characterizes the dataset, the level of complexity of the dynamics captured in it, the phenomenological features exhibited in the solution function, and the corresponding solution operator. In all these datasets, $N_{t}=21$ snapshots of each sample are collected at times $\Omega_{t}^{\Delta}=\{t_{n}=n\Delta t\}_{n=0}^{N_{t}-1}$ with uniform spacing $\Delta t$; except for the ACE and the Wave-Layer datasets where $N_{t}=20$. Each snapshot originally has a $\times128$ resolution, with uniform discretization in both spatial directions. In order to test the baselines on a random point cloud with these datasets, we randomly shuffle the coordinates and only keep the first 9216 coordinates in our experiments. Figure D.1 illustrates a few random point clouds obtained from this procedure.
Figure D.1: Random point clouds (9216 points) obtained from a $\times128$ uniform grid.
@@
Abbreviation PDE BC Domain (varies) Time Source Figure
Heat-L-Sines HE Dirichlet L-shape (✗) D.2
Wave-C-Sines WE Dirichlet circle (✗) D.3
NS-Gauss INS periodic square (✗) [13] D.4
NS-PwC INS periodic square (✗) [13] D.5
NS-SL INS periodic square (✗) [13] D.6
NS-SVS INS periodic square (✗) [13] D.7
CE-Gauss CE periodic square (✗) [13] D.8
CE-RP CE periodic square (✗) [13] D.9
ACE ACE periodic square (✗) [13] D.10
Wave-Layer WE absorbing square (✗) [13] D.11
Poisson-Gauss PE Dirichlet square (✗) [13] D.13
AF CE various airfoil (✓) [19] D.12
Elasticity FB various square with holes (✓) [19] D.14
@@ D.1.1 Incompressible Navier-Stokes equations (INS)
$\nabla\cdot\boldsymbol{v}$ $=0$ $\text{(Conservation of mass)},$ (11)
$\partial_{t}\boldsymbol{v}+(\boldsymbol{v}\cdot\nabla)\boldsymbol{v}$ $=-\nabla p+\nu\nabla^{2}\boldsymbol{v}$ $\text{(Conservation of momentum)},$
@@ D.1.5 Heat equation (HE)
D.1.6 Poisson’s equation (PE)
Poisson’s equation is given by
$\nabla u=f,$ (16)
where $f$ is a given source function and $u$ is the solution.
D.1.7 Force balance equation (FB)
@@ D.2.1 Heat-L-Sines
This dataset considers time-dependent solutions of the heat equation over an L-shaped domain which is formed by removing the rectangular portion [1, $1,\,1.5]\times[0.5$,\,1] from the larger rectangle [0.5, 1.5]×[0, $1.5]\times[0,\,$1]. We consider constant diffusivity $a=1.0$ and zero homogeneous Dirichlet boundary conditions. We prescribe a superposition of sinusoidal modes as the initial condition:
@@
where $\boldsymbol{\mu}\in[-1,1]^{d}$, $M=10$, and each $\mu_{m}$ is sampled independently and uniformly from [1,1][-1,1]. A Delaunay-based triangular mesh with characteristic length $0.008$ is used, yielding about 14,04714{,}047 nodes and 27,59027{,}590 elements. We simulate the heat equation for $100$ time steps with a time step size of $\Delta t=5\times 10^{-5}$ and keep $21$ (including the initial condition) equally spaced snapshots from each trajectory.